OpenStudy (anonymous):

convert into rectangular form r=5+5 sin theta

OpenStudy (anonymous):

as you know: \(r=\sqrt{x^2+y^2}\) and \(\theta=\arctan y/x\)

OpenStudy (anonymous):

easyest way tp proceed is just substitute this into the given equation and you got it. I am not sure if there is an easy way to represent this in rectangular coordinates.

zepdrix (zepdrix):

I was thinking you could multiply through by r,\[\Large r^2=5r+5r \sin \theta\]Making our conversion gives us something pretty ugly though... Maybe complete the square on the r's? Would that work? Hmm..\[\Large r^2-5r=5y\]

OpenStudy (anonymous):

how do you get rid of the r then?

OpenStudy (jdoe0001):

A committee is a group of people who individually can do nothing, but who, as a group, can meet and decide that nothing can be done. --- Fred Allen

OpenStudy (jdoe0001):

you can but it gets pretty jumbled then again, if you're doing polar coords, you should know what "r" equals to by now, otherwise it defeates the idea of giving you the exercise

OpenStudy (jdoe0001):

it'd be in your chapter on converting polars to rectangulars

zepdrix (zepdrix):

Did you ever figure this one out fullib? D:

OpenStudy (anonymous):

i substituted \[x^{2}+y ^{2}\] for r^2 and \[\sqrt{x^{2}+y ^2}\] for r

zepdrix (zepdrix):

Yah that seems good. I don't think you can write this in terms of \(\large y=y(x)\). I tried for a while and couldn't seem to make it work. That's prolly as good as it's gonna get. I dunno t.t

OpenStudy (anonymous):

yeah, i'm just gonna ask my teacher once school starts. thanks though!